OpenStudy (anonymous):

cot theta=-1/squareroot 3. find exact values of all theta, 0

OpenStudy (anonymous):

please help! its urgent

OpenStudy (anonymous):

\[\tan(\theta)=-\sqrt{3}\] can you solve that one?

OpenStudy (anonymous):

if you are working in degree (it looks like it) it is where sine is \(\frac{\sqrt{3}}{2}\) and cosine is \(-\frac{1}{2}\) namely at 120 degrees

OpenStudy (anonymous):

look at unit circle on last page of cheat sheet

OpenStudy (anonymous):

you will also find another one at 300 degrees, since you are looking from 0 to 360